# Word problems and embedding-obstructions in cellular automata groups on groups

### v2

@@ 8.3 The co-NEXPTIME upper bound
We now give the upper bound for the word problem that “always” holds, i.e. co-NEXPTIME.
@@
In particular, if $G$ has word problem in P (as natural groups tend to do), then it has EXPTIME word problem, but EXPTIME is strictly larger than P.
@@
We start with the first item, i.e. assume $G$ has co-NEXPTIME word problem and $X$ has PEPB with degree $t$. Let $S$ be the generating set of $G$. We show that co-word problem (nontriviality of a given element) is in NEXPTIME. Given $w$ over a generating set of automorphisms with radius $r$, guess a good $S$-labeled directed graph $H$ (which is supposed to be $B_{(nr)^{t}}$) which is equal to the $(nr)^{t}$-ball around some $h\in H$ in $H$.
For every pair of distinct elements $h,h^{\prime}\in B_{(nr)^{t}}$, check that $h^{-1}h^{\prime}$ is nontrivial in $G$. This check can be done in NEXPTIME for each individual pair (it takes at most $e^{O(((nr)^{t})^{d})}=e^{O(n^{td})}$ steps for some $d$), and since there are at most $|S|^{{nr}^{t}}$ pairs it can be done in $|S|^{(nr)^{t}}e^{O(n^{td})}=e^{O(n^{t(d+1)})}$ time in total. If the check succeeds, we do not necessarily know that $H$ is equal to the actual ball $B_{(nr)^{t}}$, but we can be sure that $B_{(nr)^{t}}$ has the graph $H$ as an image under an edge label preserving graph homomorphism (or in more group-theoretic terms, $B_{(nr)^{t}}$ covers $H$).
@@
Appendix A Complexity theory
A.1 Basic complexity theoretic definitions
@@
A.2 Polynomial space and alternating polynomial time
@@
AP is the set of problems that can be solved by alternating Turing machines in polynomial time, meaning the Turing machine can take existential transitions (all states of a nondeterministic Turing machine can be thought of as being existential, and the word is accepted at a particular step if some transition leads to an accepting computation under the stated resource bounds) and universal transitions (in which case the word is accepted if all outgoing transitions lead to an accepting computation). We omit the precise definition, which is analogous to the definition of NP (or Lemma A.3 could be taken as a definition).
@@
A.3 (A version of) Barrington’s theorem
@@
Theorem A.1 (Barrington’s theorem).
@@
Definition A.1.
Let $A,B$ be finite sets and $n\in\mathbb{N}$. Define $G_{n,A,B}$ to be the permutation group acting on $B\times A^{n}$, generated by the following permutations:
  • $\sigma$ defined by $\sigma(b,x)=(b,\sigma^{\prime}(x))$, where $\sigma^{\prime}:A^{n}\to A^{n}$ is defined by $\sigma^{\prime}(au)=ua$ where $a\in A,u\in A^{n-1}$;
  • for each even permutation $\pi$ of $B$ and $a\in A$, the map $\pi|[a]$ defined by
    $\pi|[a](b,a^{\prime}u)=\begin{cases}(\pi(b),a^{\prime}u)&\mbox{ if}a^{\prime}=a,\\ (b,a^{\prime}u)&\mbox{ otherwise}\\ \end{cases}$
One should read $\pi|[a]$ as “perform permutation‘$\pi$ given that the input is in $[a]$”, or just “$\pi$ given $[a]$”, and the idea is that we perform permutation $\pi$ on the $B$-component if the $A^{n}$-component is in the cylinder $[a]=[a]_{0}=\{au\;|\;u\in A^{n-1}\}$.
We now set up the usual calculus of controlled permutations.
First we generalize the $\pi|[a]$ notation: for any set $X\subset A^{n}$ and $\pi\in\mathrm{Alt}(B)$ (even permutations of $B$), define
@@
We refer to $A^{n}$-components of pairs $(b,u)\in B\times A^{n}$, or sets $X$ used in generators, as the control.
We will show that all these elements are in $G_{n,A,B}$ as soon as $|B|\geq 5$. For $u\in A^{*}$, define $[u]_{i}=\{w\in A^{n}\;|\;\forall j\in[0,|u|-1]:w_{i+j}=u_{j}\}$.
Lemma A.4.
For all $A,B,n$ we have $\pi|[a]_{i}\in G_{n,A,B}$.
Proof.
[…]
The generator $\sigma^{i}$ will only ever be used through this lemma. Note that $\pi|[a]_{i}$ has word norm $12i-$ 1. Thus, while technically the $A^{n}$-component is shuffled all the time in intermediate steps, one should think that the $A^{n}$-component always stays fixed and we are simply conditioning permutations $\pi$ on different positions. The reason we explicitly use the generator $\sigma$ is to have a uniform generating set, and so that $\pi|[a]_{i}$ has word norm $i$ (as it will in the applications).
Lemma A.5.
@@
Remark A.1.
The statement of this lemma is nice and clean, but a few remarks on its necessity may be in order, since we are not aware of a proof that avoids case distinctions. The lemma was stated by Ore in [24], and proved by Ito in [20]. Ito’s paper seems difficult to find, but a proof (of a stronger result about even permutations) is found also in [5]. It is known that the commutator width is one in all simple groups. We need the lemma only for a fixed set of cardinalities $|B|$, so one could simply check it by computer. It even suffices that the commutator width if bounded (over all $|B|$), although this changes the degrees of polynomial bounds. One can even phrase the argument directly in terms of $3$-rotations to avoid using the lemma entirely by using simply $[(a\;b\;c),(c\;d\;e)]=(c,b,e)$; this latter is the approach taken by Barrington. We also sometimes use only $3$-rotations, namely in some cases we need the support to be non-full.
Lemma A.8.
@@
We note that for complexity-theoretic purposes, this is an inefficient definition, and one should balance the brackets instead. However, we only do this for commutators of bounded length, and this simple definition sufficies.
Lemma A.10.
@@
Theorem A.2 (variant of Barrington’s theorem).
@@
Let $A=\{0,1\}$. Let $B$ be any alphabet with cardinality at least $5$. Let $s=(s_{1},s_{2}):\mathrm{Alt}(B)\to\mathrm{Alt}(B)\times\mathrm{Alt}(B)$ be any function such that $\pi=[s_{1}(\pi),s_{2}(\pi)$ for all $\pi$, which exists by Lemma A.7. Identify Boolean formulas $\phi$ with the subsets of $A^{n}$ they describe. Then by Lemma A.8 and Lemma A.10 we have
@@
Let $T_{d}$ be the maximal word norm of $\pi|\phi$ for $\phi$ of depth $k=O(\log n)$, when using generators of the form $\pi|[a]_{i}$ for $a\in A$. The formulas show $T_{i}\leq 6T_{i-1}$ for all $i$, so $T_{k}=O(6^{k})=O(6^{\log n})=O(n^{d})$ for some $d$. Finally, each $\pi|[a]_{i}$ has word norm $O(n)$. ∎
We now give the main examples of languages are in BWBP’.
@@
By Theorem A.2, $\pi|L$ has polynomial norm in $G_{n,A,B}$ for $|B|\geq 5$. ∎
@@
Lemma A.14.
The language $\{\mathrm{unary}_{m}(n)\#\mathrm{unary}_{m}(n+1)\;|\;n\in[0,m)\}\subset\{0,1,\#\}^{*}$ is in BWBP’. Similarly, the relation $\{u2^{n}\#ua2^{n}\;|\;u\in\{0,1\}^{*},n\in\mathbb{N}\}$ is in BWBP’ for each $a\in\{0,1\}$.
@@
For simplicity assume $m$ is a power of $2$. For the divide-and-conquer, observe that checking the input to be in the form $\mathrm{unary}_{m}(n)\#\mathrm{unary}_{m}(n+1)$ can be split in two cases after splitting the input in half:
@@
The other relation is shown to be in BWBP’ by first checking that the projection $0,1\mapsto 0;2\mapsto 1$ give a word in the first language, then checking the binary symbols are in the same position, or in case only the second word carries a bit, this bit is $a$. (This latter is a matter of checking that all pairs of corresponding positions in the two parts belong to a particular set.) ∎
@@
Lemma A.15.
The language $\{\mathrm{bin}_{m}(n)\#\mathrm{bin}_{m}(n+1)\;|\;n\in[0,2^{m})\}\subset\{0,1,\#\}^{*}$ is in BWBP’.
@@
Lemma A.16.
Fix $d$, $|B|\geq 5$ and any $\pi\in\mathrm{Alt}(B)$. Then for any $\ell$, the permutation $\pi|\{\mathrm{bin}_{m}(n_{1})\#\mathrm{bin}_{m}(n_{2})\#\cdots\#\mathrm{bin}_{m}(n_{d})\;|\;\sum_{i}n_{i}=\ell\}$ has word norm polynomial in $m$ in $\mathcal{G}_{dm+d-1,A,B}$.
@@
It suffices to construct a circuit whose output is the sum of two numbers, as one can chain $d-1$ of these together and finally compare the sum to a fixed number using Lemma A.13.
This is [1, Example 6.19]. The bits of the sum are determined by the added bits at that position, as well as the carry signal coming from the left (from the left, since we use LSB for binary numbers). It is easy to show that the carry signal at a particular position can in turn be “predicted” by a constant-depth circuit, when we allow conjunctions and disjunctions (circuit nodes labeled $\wedge$ and $\vee$) with arbitrarily many inputs. Associating these in a balanced way gives a logarithmic depth circuit where all nodes have in-degree $2$. ∎
Remark A.2.
In modern CPU design this corresponds roughly to the carry-lookahead adder construction (speeding up the school algorithm of linear carry propagation, which in turn is known as a ripple-carry adder in circuit design); Babbage invented a similar anticipating carriage for his analytical engine.

### v3

@@ 9.3 The co-NEXPTIME upper bound
We now give the upper bound for the word problem that “always” holds; of course, we need some technical assumptions on the complexity of the group $G$ itself, and the complexity of the subshift.
@@
In particular, if $G$ has word problem in P (as natural groups tend to do), then it has EXPTIME word problem. EXPTIME is known to be strictly larger than P, and there are also groups with word problem in EXPTIME $\setminus$ P.
@@
We start with the first item, i.e. assume $G$ has co-NEXPTIME word problem and $X$ has PEPB with degree $t$. Let $S$ be the generating set of $G$. We show that the co-word problem (nontriviality of a given element) is in NEXPTIME. Given $w$ over a generating set of automorphisms with radius $r$, guess a good $S$-labeled directed graph $H$ (which is supposed to be the actual ball $B_{(nr)^{t}}(1)$ in $G$). Check (deterministically) that it is equal to the $(nr)^{t}$-ball $B_{(nr)^{t}}(h_{0})$ around some $h_{0}\in H$ in $H$.
For every pair of distinct elements $h,h^{\prime}\in B_{(nr)^{t}}(h_{0})$ in the graph $H$, check that $h^{-1}h^{\prime}$ is nontrivial in $G$. This check can be done in NEXPTIME for each individual pair (it takes at most $e^{O(((nr)^{t})^{d})}=e^{O(n^{td})}$ steps for some $d$), and since there are at most $|S|^{{nr}^{t}}$ pairs it can be done in $|S|^{(nr)^{t}}e^{O(n^{td})}=e^{O(n^{t(d+1)})}$ time in total. If the check succeeds, we do not necessarily know that $H$ is equal to the actual ball $B_{(nr)^{t}}$, but we can be sure that $B_{(nr)^{t}}$ has the graph $H$ as an image under an edge label preserving graph homomorphism (or in more group-theoretic terms, $B_{(nr)^{t}}$ covers $H$).
@@
10 Open questions
While our approach shows that $A^{\mathbb{Z}^{D}}$ does not embed in $B^{\mathbb{Z}^{d}}$ for $D>d$, it seems not to be helpful for comparing $A^{\mathbb{Z}^{d}}$ and $B^{\mathbb{Z}^{d}}$ for distinct alphabets. (This is another embeddability question of Hochman.)
Also, while we obtain several nonembeddability results (and we ask several technical results concerning the extensions of these results), the following general problem stays open:
Question 10.1.
Are there nontrivial embeddings of $\mathrm{Aut}(A^{G})$ into $\mathrm{Aut}(B^{H})$, for some nontrivial alphabets $A,B$, and f.g. infinite groups $G,H$, such that $G$ is not a free group?
Of course, embeddings exist when $G$ is a subgroup of $H$, and $|A|\leq|B|$. However, we do not know any other examples.
We showed in Theorem 8.1 that the automorphism group of a full shift on the lamplighter group has co-NEXPTIME complete word problem, which as we showed in Theorem 9.12 is in a sense the maximal possible. It is an interesting question whether we can reach this upper bound in a large class of groups.
Question 10.2.
For which groups $G$ does $\mathrm{Aut}(A^{G})$ have co-NEXPTIME-complete word problem?
If one can Lipschitz embed a copy of the lampligher group in a group $G$, then one should be able to obtain the hardness result with the usual method of Section 6, at least with some alphabets. However, at present we do not know how common this is.
In any case, it seems likely that this method is suboptimal, and one should instead consider a generalized notion of “biarboreousness” (the lamplighter group is a horocyclic product of trees, and this is precisely what allows us to perform computation on an exponential sized grid), allowing for contracting paths instead of strict trees. But again, we do not know how commonly such structures are found in groups.
There are many questions left about simplistic growth. We showed in Theorem 5.1 that for groups with stretched exponential growth, one has a certain lower bound (far surpassing the corresponding upper bound on free abelian groups) on simplistic growth. This shows that such groups do not embed in $\mathrm{Aut}(A^{\mathbb{Z}^{d}})$ (or other groups with polynomial systolic growth). However, we do not immediately obtain nonembeddability results between other $\mathrm{Aut}(A^{G})$ groups, among groups $G$ of stretched exponential growth.
By optimizing the degree $\ell$, which relates to both the growth rate of the group and the efficiency of implementation of the scheme, we would possibly allow proving many more noninclusions between $\mathrm{Aut}(A^{G})$ groups. However, at present it is not clear what the optimal approach is.
Question 10.3.
Can we in general attain simplistic growth in $\mathrm{Aut}(A^{G})$, which is closely driven by the growth $f$ of the group $G$? For example, can we always (or in a large class of groups) reach $e^{f(n)/\mathrm{poly}\log(n)}!$ simplistic growth?
In the case of groups of polynomial growth (equivalently, virtually nilpotent groups), it seems likely that one can use normal forms for group elements, and implement a variant of Lemma 4.4 to always attain roughly the (factorial of the exponential of the) growth rate, as we do in the case of abelian groups.
We note, however, that this approach is fundamentally limited, and the maximal simplistic growths we can attain with such an approach are of the form $e^{2^{\sqrt[d]{n}}}!$. It is certainly possible to have much higher simplistic growths in some groups. For example, if the group $G$ contains an element with prime order $p$ and word norm $n$, then the corresponding right shift shows that the simplistic growth satisfies $f(n)=p$.
There is in fact no limit to how quickly elements with high (but finite) prime order can be reached in $2$-generated groups, by (the proof of) the embedding theorem in [20]. Specifically, this proof shows that for each $n$, we can freely choose the order of a single element with word norm $O(n)$. In the residually finite case, we have an upper bound on simplistic growth in terms of the systolic growth, so it would be natural to try to show that the systolic growth (rather than the growth of balls) can be reached, or at least approached, in a large class of groups.
Question 10.4.
Can we in general attain simplistic growth in $\mathrm{Aut}(A^{G})$, which is closely driven by the systolic growth of the group $G$?
Finally, the technique of Section 9.2 is based on “polynomial splitting schemes”, whose existence we show for free and surface groups (and slightly beyond).
Question 10.5.
Which groups admit polynomial splitting schemes?
We leave it as an open question what groups admit such schemes. One interesting question is whether they exist for small cancellation groups or hyperbolic groups with one-dimensional boundary.
Appendix A Index
$A$ an alphabet, often $B\times C$ in case of $\mathrm{PAut}$
$\mathrm{Aut}(X)$ automorphism group of a subshift $X$
$B$ the control part of an alphabet (for controlled permutations), or coordinate part (for ripple catching)
$\beta_{\pi,\ell}$ applies $\pi$ on the $C$-component of the color of a node $u$ if the $B$-component of $u$ contains $b$ such that $\rho(b)=\ell$.
$C$ the permuted part of an alphabet
$d,D$ dimensions; or $d$ can also be a distance
$\eta_{\pi}\in\mathrm{PAut}(A)$ Symbol permutation $\pi\in\mathrm{Sym}(A)$
$f,g$ functions
$G$ the shift group, acting on $A^{G}$
$\gamma_{\pi}$ apply $\pi$ at a node (in the $C$-component) if that node is successful
$\mathcal{G},\mathcal{H}$ a subgroup of $\mathrm{Aut}(X)$
$\mathcal{G}_{m,C,B}$ the group of controlled permutations
$\mathcal{G}_{S,C,B}$ the ripple group
$i,j,k,m,n,p$ numbers, various uses
$\mathcal{Q}_{S,C,B}$ good $S$-labeled $B\times C$ colored graphs
$\mathrm{PAut}(A),\mathrm{PAut}[B;C]$ groups generated by partial shifts and symbol permutations
$\pi$ permutation, usually $\pi\in\mathrm{Sym}(C)$
$\pi_{b,i}\in\mathcal{G}_{m,C,B}$ the generator that applies $\pi$ if the $i$th bit of the $B$ component contains $b$
$\phi_{\pi,c}\in\mathcal{G}_{S,C,B}$ apply $\pi$ in $C$-component of node $u$ if $S$-neighbors of $u$ (that exist) have symbol $c$
$\phi_{\pi,c,s}\in\mathrm{PAut}(A)$ $\phi_{\pi,c}$-analog (for a single $s\in S$) in automorphism group
$\phi_{\pi,c,S}\in\mathrm{PAut}(A)$ $\phi_{\pi,c}$-analog in automorphism group
$\psi_{\pi,\ell,c}$ apply $\pi$ in $C$-component of node $u$ if rank of $u$ is $\ell$ and $S$-neighbors of $u$ have symbol $c$
$\rho$ rank function of a good ranked graph
$\sigma_{g,i}$ partial (right!) shift by $g$ on track $i$
$s$ edge-label, or generator of a free group
$S$ finite set of edge-labels, or generators of a free group; often acts on labels
$t$ word over edge-labels
$u,v$ nodes of an abstract graph; or finite words
$\vec{u},\vec{v}$ vectors in $\mathbb{Z}^{d}$ or $\mathbb{Z}^{D}$
$w$ finite word
$x,y,z$ configurations in a subshifts
Appendix B Complexity theory
B.1 Basic complexity theoretic definitions
@@
B.2 Polynomial space and alternating polynomial time
@@
AP is the set of problems that can be solved by alternating Turing machines in polynomial time, meaning the Turing machine can take existential transitions (all states of a nondeterministic Turing machine can be thought of as being existential, and the word is accepted at a particular step if some transition leads to an accepting computation under the stated resource bounds) and universal transitions (in which case the word is accepted if all outgoing transitions lead to an accepting computation). We omit the precise definition, which is analogous to the definition of NP (or Lemma B.4 could be taken as a definition).
@@
B.3 (A version of) Barrington’s theorem
@@
Theorem B.5 (Barrington’s theorem).
@@
Definition B.6.
Let $C,B$ be finite sets and $n\in\mathbb{N}$. Define $G_{n,C,B}$ to be the permutation group acting on $C\times B^{n}$, generated by the following permutations: for each permutation $\pi$ of $C$ and $b\in B$ and $i\in\{0,\ldots,n-1\}$, the map $\pi_{b,i}$ defined for $|u|=i,b\in B,|v|=n-i-1,c\in C$ by
$\pi_{b,i}(c,ub^{\prime}v)=\begin{cases}(\pi(c),ub^{\prime}v)&\mbox{ if}b^{\prime}=b,\\ (b,ub^{\prime}v)&\mbox{ otherwise}\\ \end{cases}$
We endow $G_{n,A,B}$ with the word norm coming from these generators.
In English, $\pi_{a,i}$ performs permutation $\pi$ on the $B$-component if the $A^{n}$-component has symbol $a$ at $i$, and we read it as “$\pi$ controlled by $a$ at $i$”.
We set up a calculus of controlled permutations.
For any set $X\subset A^{n}$ and $\pi\in\mathrm{Alt}(B)$ (even permutations of $B$), define
@@
We refer to $A^{n}$-components $u$ of pairs $(b,u)\in B\times A^{n}$, or sets $X$ used in generators, as the control. Note that $\pi_{a,i}=\pi|[a]_{i}$ where $[a]_{i}$ is the cylinder $[a]_{i}=\{uav\;|\;u\in A^{i},v\in A^{n-i-1}\}$.
Lemma B.7.
@@
Remark B.10.
The statement of this lemma is nice and clean, but a few remarks on its necessity may be in order, since we are not aware of a proof that avoids case distinctions. The lemma was stated by Ore in [25], and proved by Ito in [22]. Ito’s paper seems difficult to find, but a proof (of a stronger result about even permutations) is found also in [5]. It is even known that an analogous result holds (i.e. the commutator width is one) in all simple groups. We need the lemma only for a fixed set of cardinalities $|B|$, so one can easily give an ad hoc proof. It even suffices that the commutator width if bounded (over all $|B|$); this changes the degrees of polynomial bounds but none of our main statements. One can also phrase all our arguments directly in terms of $3$-rotations to avoid using the lemma entirely by using simply $[(a\;b\;c),(c\;d\;e)]=(c,b,e)$; indeed this is a common way to prove Barrington’s theorem. We also sometimes use only $3$-rotations in the present paper, namely in some cases we need the support to be non-full.
Lemma B.11.
@@
For complexity-theoretic purposes, this is an inefficient definition, and one should balance the brackets instead. However, we only do this for commutators of bounded length, and this simple definition sufficies.
Lemma B.13.
@@
Theorem B.16 (variant of Barrington’s theorem).
@@
Let $A=\{0,1\}$. Let $B$ be any alphabet with cardinality at least $5$. Let $s=(s_{1},s_{2}):\mathrm{Alt}(B)\to\mathrm{Alt}(B)\times\mathrm{Alt}(B)$ be any function such that $\pi=[s_{1}(\pi),s_{2}(\pi)$ for all $\pi$, which exists by Lemma B.9. Identify Boolean formulas $\phi$ with the subsets of $A^{n}$ they describe. Then by Lemma B.11 and Lemma B.13 we have
@@
Let $T_{d}$ be the maximal word norm of $\pi|\phi$ for $\phi$ of depth $k=O(\log n)$. The formulas show $T_{i}\leq 6T_{i-1}$ for all $i$, so $T_{k}=O(6^{k})=O(6^{\log n})=O(n^{d})$ for some $d$. ∎
We now give the main examples of languages are in BWBP’, by showing they are in NC1.
@@
By Theorem B.16, $\pi|L$ has polynomial norm in $G_{n,A,B}$ for $|B|\geq 5$. ∎
@@
Lemma B.19.
The language $\{\mathrm{unary}_{m}(n)\#\mathrm{unary}_{m}(n+1)\;|\;n\in[0,m)\}\subset\{0,1,\#\}^{*}$ is in BWBP’. Similarly, the relation $\{u2^{n}\#ua2^{n-1}\;|\;u\in\{0,1\}^{*},n\in\mathbb{N}\}$ is in BWBP’ for each $a\in\{0,1\}$.
@@
First, consider th first language. For simplicity assume $m$ is a power of $2$. For the divide-and-conquer, observe that checking the input to be in the form $\mathrm{unary}_{m}(n)\#\mathrm{unary}_{m}(n+1)$ can be split in two cases after splitting the input in half:
@@
Consider now the second language. It is in BWBP’, as it suffices to first check that the projection $0,1\mapsto 0;2\mapsto 1$ give a word in the first language, then check that binary symbols appear in the same positions in the words on each side of ##\##, or in case only the second word carries a bit, this bit is $a$. (All but the first check are a matter of checking that all pairs of corresponding positions in the two parts belong to a particular set.) ∎
@@
Lemma B.20.
The language $\{\mathrm{bin}_{m}(n)\mathrm{bin}_{m}(n+1)\;|\;n\in[0,2^{m})\}\subset\{0,1\}^{*}$ is in BWBP’.
@@
Lemma B.21.
Fix $d$, $|C|\geq 5$ and any $\pi\in\mathrm{Alt}(C)$. Then for any $\ell$, the permutation $\pi|\{\mathrm{bin}_{m}(n_{1})\mathrm{bin}_{m}(n_{2})\cdots\mathrm{bin}_{m}(n_{d})\;|\;\sum_{i}n_{i}=\ell\}$ has word norm polynomial in $m$ in $\mathcal{G}_{dm+d-1,C,B}$.
@@
It suffices to construct a circuit whose output is the sum of two numbers, as one can chain $d-1$ of these together and finally compare the sum to a fixed number using Lemma B.18.
This is [1, Example 6.19]. The bits of the sum are determined by the added bits at that position, as well as the carry signal coming from the left (from the left, since binary numbers are written least significant bit first). It is easy to show that the carry signal at a particular position can in turn be “predicted” by a constant-depth circuit, when we allow conjunctions and disjunctions (circuit nodes labeled $\wedge$ and $\vee$) with arbitrarily many inputs. Associating these in a balanced way gives a logarithmic depth circuit where all nodes have in-degree $2$. ∎